If Maria estefan disposable IncomeIncreaseFrom 600 to 650

And her level of personal consumption expenditures increase from 450 to 520 You may conclude that’s her marginal propensity to


1
Expert's answer
2022-04-13T09:19:20-0400

"i=\frac{\frac{650-600}{600}}{\frac{520-450}{450}}=0.54."